# SceneSplat: Gaussian Splatting-based Scene Understanding with Vision-Language Pretraining

### v1

@@ F Limitations
Our dataset quality is largely influenced by the source indoor datasets. Low resolution and blurry images can cause floating artifacts. Downstream tasks are also limited by the original dataset’s annotations. Temporal inconsistency in the current language label obtaining process needs to be addressed as it can pollute the vision-language pretraining. Also, we plan to add the bounding boxes [13] and language descriptions [27] as next step.
G Impact Statement
@@
This work introduces SceneSplat-7K, the first large-scale dataset of 3D Gaussian Splatting (3DGS) specifically designed for indoor scene understanding. By providing over 7,000 annotated scenes, it enables standardized benchmarking for 3DGS-based reasoning in vision tasks, paving the way for more robust 3D scene comprehension. The proposed framework for vision-language pretraining tailored to 3DGS enhances semantic alignment by incorporating language-annotated 3D Gaussian splats, which establish clear connections between latent representations and 3D scenes. Leveraging knowledge distillation from 2D foundation models, our approach pushes the boundaries of 3D vision-language understanding, achieving state-of-the-art performance in open-vocabulary tasks. Furthermore, our model, combining contrastive learning and masked point modeling (MPM), significantly outperforms existing methods in 3D semantic segmentation. This work lays a critical foundation for advancing scalable 3D scene understanding, with broad implications for autonomous systems, robotics, and augmented reality applications. By releasing the dataset, model, and code publicly, we foster further innovation and facilitate the development of generalizable solutions for 3D vision tasks.

### v2

@@ E Limitations
The quality of our dataset is largely influenced by the source indoor datasets. Low-resolution and blurry images can cause floating artifacts. Downstream tasks are also limited by the annotations of the original dataset. Temporal inconsistency in the current language label obtaining process needs to be addressed, as it can pollute vision-language pretraining. In addition, we plan to add bounding boxes [13] and language descriptions [27] in the next step.
F Impact Statement
@@
This work introduces SceneSplat-7K, the first large-scale indoor dataset of 3D Gaussian Splatting (3DGS). By providing over 7K annotated scenes, it enables standardized benchmarking for 3DGS-based reasoning in vision tasks. The proposed framework for vision-language pretraining tailored to 3DGS enhances semantic alignment and establishes a clear connection between latent representations and 3DGS scenes. By leveraging knowledge distillation from 2D foundation models, combining contrastive learning and masked gaussian modeling (MGM), our approach significantly outperforms existing methods in 3D semantic segmentation. This study lays the foundation for advancing scalable 3D scene understanding, with broad implications for autonomous systems and augmented reality applications. By publicly releasing the dataset, model, and code, we foster further innovation and facilitate the development of generalizable solutions for 3D scene understanding.